# Lan-grasp: Using Large Language Models for Semantic Object Grasping

### v2

@@ 4.7 Conclusion and Future Work
In this paper, we presented Lan-grasp, a novel approach to semantic object grasping. By leveraging foundation models, we provide our approach with a deep understanding of the objects and their intended use in a zero-shot manner. Through extensive experiments, we showed that for a wide range of objects Lan-grasp is generating grasps that are preferred by humans and also ensure safety and object usability. In particular, the proposed grasps were compared to human preferences gathered through a questionnaire. The evaluations showed that Lan-grasp performs consistently better on that metric than the baseline methods. We also proposed a feedback loop approach that reasons about grasp feasibility in complex scenarios. In future work, we plan to test Lan-grasp in more complex, and cluttered environments to evaluate its robustness. Additionally, we aim to enhance the feedback loop by introducing mechanisms for when no feasible grasp is detected. For instance, the robot could ask a human for assistance or employ more sophisticated reasoning strategies to modify the environment to facilitate grasping. This would make our algorithm capable of handling more complex real-world scenarios. Inspired by these results, in the future we plan to further exploit Large Language Models to not only decide where to grasp an object but also how to grasp and hold it according to a specific task. As an example, we would expect a robot operating in daily environments to hold a knife vertically and downwards when the task is to carry the knife around rather than holding the knife in a horizontal pose. This would be the next step towards more meaningful grasps that help robots with object manipulation and task execution in day-to-day environments.

### v3

@@ 4.8 Conclusion and Future Work
In this paper, we presented Lan-grasp, a novel approach to semantic object grasping. By leveraging foundation models, we provide our approach with a deep understanding of the objects and their intended use in a zero-shot manner. Through extensive experiments, we showed that for a wide range of objects Lan-grasp is generating grasps that are preferred by humans and also ensure safety and object usability. In particular, the proposed grasps were compared to human preferences gathered through a questionnaire. The evaluations showed that Lan-grasp performs consistently better on that metric than the baseline methods. We also proposed a feedback loop approach that reasons about grasp feasibility in complex scenarios. In future work, we plan to test Lan-grasp in more complex and cluttered environments to evaluate its robustness in generating meaningful grasps. In such settings, one possible approach is to first segment the object and composite it onto a plain background before prompting the VLM to localize the desired object part, similar to preprocessing used in prior visuomotor policy work (Mirjalili et al., 2025). Additionally, we aim to enhance the feedback loop by introducing mechanisms for when no feasible grasp is detected. For instance, the robot could ask a human for assistance or employ more sophisticated reasoning strategies to modify the environment to facilitate grasping. This would make our algorithm capable of handling more complex real-world scenarios. Inspired by these results, in the future we plan to further exploit Large Language Models to not only decide where to grasp an object but also how to grasp and hold it according to a specific task. As an example, we would expect a robot operating in daily environments to hold a knife vertically and downwards when the task is to carry the knife around rather than holding the knife in a horizontal pose. This would be the next step towards more meaningful grasps that help robots with object manipulation and task execution in day-to-day environments.
We also presented a method for placing a grasped object in an upright position using large pretrained models. Despite high success rates on some items, the object pose estimate from SAM 3D is not always sufficiently precise, especially when the object is turned away from the camera or only the object’s bottom part is visible. We currently tackle this issue by rotating the gripper by $90^{\circ}$, but in future a more principled approach could be implemented to ensure optimal object visibility. Alternatively, pose estimation could be improved using explicit models for 3D-to-2D matching in conjuncture with SAM 3D meshes, e.g., FoundationPose (Wen et al., 2024).